OpenStudy (brooklynriojas1):

find the missing side or angle of the triangle. Round to the nearest tenth. C(angle)=53 degrees B(angle)= 80 degrees a(side)=2 b(side)=?

OpenStudy (jackthegreatest):

i believe you find angle a

OpenStudy (brooklynriojas1):

How do I find angle a with the information given?

OpenStudy (jackthegreatest):

180-53-80 = angle a

OpenStudy (brooklynriojas1):

oh you right, thank you.

OpenStudy (jackthegreatest):

\[\frac{ Side A }{ Sin (Angle A)}=\frac{ Side B}{ Sin (Angle B)}\]

OpenStudy (jackthegreatest):

there is the correct equation

OpenStudy (anglina4196):

180-(53+80)